# Statistical modeling of SOFR term structure

### v2

@@ 1 Introduction
Due to the illiquidity and incompleteness, classical arbitrage arguments are not well suited for pricing SOFR derivatives. Indeed, risk-neutral expectations yield linear pricing rules that are poor descriptions of observed market prices or the prices at which an agent might be willing to trade. In the face of incompleteness, the replication arguments behind risk-neutral valuations do not hold. In incomplete and illiquid markets, offered prices are best described by the indifference pricing principle, which is based on approximate hedging arguments capturing nonlinearities and the subjectivity of offered prices. Indifference pricing builds on a subjective (as opposed to risk-neutral) probabilistic description of the underlying rates, prices, and other relevant risk factors such as inflation and GDP; see e.g. [car9, pen14].
@@
The model presented in this paper captures all the above features. The model is easy to calibrate to historical data, user’s views and observed futures rates. It is fast in simulations, which makes it well-suited for both pricing and risk management applications. We illustrate the simulation model with indifference pricing of SOFR futures, options and swaptions. The pricing exercises here are meant only as an illustration of the convenience of the model in large-scale simulations. The pricing and hedging applications will be developed further in a follow-up paper, that builds on computational optimization of semi-static hedging strategies employing hundreds of exchange-traded SOFR derivatives.
@@
Term structure models under $P$-measure are widely used by public organisations and asset managers. Many of them are based on the Nelson-Siegel and Svensson parameterizations, which describe the yield curve with three or four parameters, respectively; see [nelson1987parsimonious, svensson1994estimating]. For example, [diebold2006macroeconomy] develops a stochastic term structure model where the yield curve is statistically connected to macroeconomic variables. An extension with stochastic volatility was proposed in [bianchi2009great] while [yu2011forecasting] incorporated corporate bonds. The model presented in this paper allows for more flexible parameterizations, and it incorporates jumps on FOMC meeting dates, which is an essential feature of the overnight lending market.
@@
In derivative pricing, term structure models are often specified so that the corresponding zero-coupon bond prices have arbitrage-free dynamics. A common approach is to specify the term structure dynamics together with a risk-neutral measure (or, more generally, a pricing kernel) under which zero-coupon bond prices of all maturities are martingales. Risk-neutral term structure models for SOFR have been developed in [bickersteth2021pricing, brace2024sofr, fang2024drives, fontana2024term, gellert2021short, lyashenko2019looking, schlogl2024term, skov2021dynamic] and their references. Statistical analysis in [fang2024drives] finds that jumps in SOFR are driven by macroeconomic factors, but their stochastic term structure model does not capture the effect. Closer to the present work are the discrete-time models of [ang2003no, lemke2008affine] which describe the joint dynamics of the term structure and relevant macroeconomic variables. In incomplete markets, the martingale measure (or the pricing kernel) is not unique and the chosen one is not necessarily the one that best captures statistical features of the term structure. Like our model, those of [ang2003no] and [lemke2008affine] describe the influence of inflation and GDP growth on the central bank policy rate but, contrary to the main role of central banking, however, the model of [ang2003no] does not describe the influence of interest rates on the macroeconomy. Moreover, the models of both [ang2003no] and [lemke2008affine] involve latent variables, which makes the calibration of the models complicated. Our model, on the other hand is very easy to calibrate to both historical data, current term structure and user’s views. Neither [ang2003no] nor [lemke2008affine] describe the possible jumps on the FOMC meeting dates.
@@ 2.1 SOFR averages and futures
For a three-month SOFR futures contract, the reference period $t_{0},t_{1}]$ starts on the third Wednesday of the third calendar month before the delivery month and ends on the third Wednesday of the delivery month. According to the CME Rulebook [CMESOFRFut3m, Section 46003.A], the payout of a long position is delivered at time $t_{1}$ and given by
@@
where $F(t_{0},t_{1})$ is the futures rate agreed between the counterparties. The exchange-traded three-month futures rates come with bid-ask spreads. We will denote the bid- and ask-rates by $F^{b}(t_{0},t_{1})$ and $F^{a}(t_{0},t_{1})$, respectively. If we want to emphasize that the quotes were observed at time $t$, we will add a subindex and write $F_{t}^{b}(t_{0},t_{1})$ and $F_{t}^{a}(t_{0},t_{1})$.
@@
At CME, SOFR futures are quoted in terms of the “futures price” $f_{t}(t_{0},t_{1})$ which is related to the futures rate above by $F_{t}(t_{0},t_{1})=1-f_{t}(t_{0},t_{1})/100$; see [imm].
@@
where $F(t_{0},t_{1})$ is the futures rate agreed between the couterparties.
@@ 3 Parameterization of the forward curve
In practice, the term structure of interest rates is relevant only up to a certain cut-off date $\bar{T}$. For example, if one wishes to estimate the forward curve five years into the future, we would have $\bar{T}=5\times 365$. The number of futures and swap quotes in practice is not enough to uniquely fix the whole forward curve $(F(t))_{t=0}^{\bar{T}}$ through condition (8). Following [fil9, Section 9.3], we will approximate the overnight forward curve by an affine model of the form
@@
Note that equation (10) is linear in the parameters $\xi=(\xi^{k})_{k=1}^{K}$ so the above least squares problem is quadratic in the parameters $\xi$. Besides futures quotes, one could use quotes on zero-coupon bonds in the calibration. Indeed, zero-coupon bonds imply futures rates $F(t_{0},t_{1})$ through (3) so the corresponding consistency conditions would also be linear. Swap contracts, however, would result in nonlinear consistency conditions for the forward rates so their inclusion in the least squares estimation would result in a non-convex optimisation problem. Such problems are NP-hard, in general; see e.g. [nes18].
@@
Problem (14) is that of convex quadratic optimization so it can be solved quickly by off-the-shelf optimization solvers. In this paper, we employ the interior point solver of Mosek [aps2019mosek]. The problem was formulated and communicated to Mosek using Python [van1995python] and CVXPY [diamond2016cvxpy]. The solution is obtained in less than 0.4 seconds using a Dell Latitude running Linux with 16 GB of RAM and a 4-core intel i7 CPU at 3.00 GHZ; see Example 4. Similarily, solving 1300 instances of problem (11) takes less than 8 minutes; see Section 4.2.
@@
Such basis functions result in piecewise constant forward curves which are used e.g. by CME [cme]. Like CME, we use one-month and three-month futures quotes and SOFR to estimate the forward curve. However, instead of mid-quotes, we use the bid and ask quotes, and solve the least squares problem (14) with tenors corresponding to FOMC meeting dates
@@
$\phi^{k}(t)=\begin{cases}1&t=T_{k},\\ 0&t\notin[T_{k-1},T_{k+1}].\end{cases}$
This results in continuous piecewise-linear forward curves. Indeed, the corresponding forward curve given by (9), has $F(T_{k})=\xi^{k}$ for all $k$ and the curve interpolates linearly between consecutive tenors. Continuity is often seen as a desirable property in term structure modelling; see e.g. [fil9, Section 3.3] and the references there. Having an economic interpretation for the parameters $\xi^{k}$, facilitates their statistical modelling; see Section 4.
@@
In particular, the first component of $\xi$ is the one-day maturity zero-rate $F(0)$ which is essentially the current SOFR; see (7). We fit the above parameterization to the observed futures quotes as well as to the value of SOFR observed the morning after. We do this by by solving problem (14) with the additional constraint that $\xi^{1}=\ln(1+F(0,1)\delta)/\delta$; see equation (7). This results in the piecewise linear curve in Figure 2. The figure also plots the employed futures bid- and ask-quotes obtained from Bloomberg Finance L.P. Because the SOFR rate is published on the next business day, we shift the horizontal position of the quotes 1 day backwards to match with the futures prices that are published on the trading day.
@@ 4 Statistical modelling of the forward curve
Up to now, we have been studying the term structure at a single point in time. We will next focus on the dynamics of the term structure and denote the forward curve at time $t$ by $F_{t}:=(F_{t}(s))_{s=0}^{T}$. In order to describe the evolution of $F_{t}$, we will, much like in Chapter 9 of [fil9], first approximate it by the parameterization (9) and then model the parameters as a multivariate stochastic process. More precisely, we will assume that $F_{t}$ has, for every day $t=0,1,2,\ldots$, the parametric representation
$F_{t}(s)=\sum_{k=1}^{K}\xi^{k}_{t}\phi^{k}(s-t)\quad s\geq t,$
where $\xi^{k}:=(\xi^{k}_{t})_{t=0}^{\infty}$ are real-valued stochastic processes. We will denote by $\xi$ the $\mathbbm{R}^{K}$-valued stochastic process with components $\xi^{k}$. This can be seen as a parametric term structure model in discrete time much as e.g. in [fil9, Section 9.3]. While term-structure models are usually specified under a fixed risk-neutral measure, we will describe the behavior of the term structure under a “statistical” measure (also known as the $P$-measure” or the “real-world measure”) which will be more suitable for portfolio optimization and indifference pricing of SOFR derivatives. We will follow the general procedure described in [aap21, Sections 3 and 4] to develop a stochastic model for $\xi$. This allows us to calibrate the model to both historical data and user’s views concerning the future development of the term structure.
@@ 4.1 SOFR and policy rates
SOFR is calculated daily as a volume-weighted median rate of overnight repurchase transactions (repo) of depositary institutions and the FED; see [sifmarepo]. The FED offers a deposit rate, known as overnight Reverse Repurchase Agreement Facility (ON RRP) since September 2014, and a lending rate, known as Standing Overnight Repurchase Agreement Facility (SRF) since July 2021; see Table 1. Most participants in the repo market have access to these Fed facilities (see [RRepoCP, SRepoCP, secrepo]) so they are unlikely to lend to a counterparty at a rate lower than ON RRP. Similar to the federal funds target range $[L,H]$, SRF and ON RRP are set during Federal Open Market Committee (FOMC) meetings and remain constant between meetings. Historically, ON RRP was set above $L$, except for a brief period between the end of 2019 and the beginning of 2020; see Figure 3. When enough liquidity exists in the FED reverse repo facility, one would expect SOFR to stay above $L$ as well. The rare cases where SOFR falls below $L$ may be explained by the demand for ON RRP exceeding what the FED is offering, or some participants, which do not have access to FED facilities, accepting deposit rates lower than ON RRP. The converse applies to the federal funds target range upper limit $H$. Table 1 illustrates the situation on 4 January 2024.
@@
Table 1: A glimpse of FED monetary policy measures as of 4 January 2024. SOFR was $5.32\%$ (source: FED [FederalReserveWebsite].)
@@
Figure 3: Daily values of SOFR, ON RRP, and the limits of federal funds target range $[L,H]$ (source: FRED [SOFR, DFEDTARU, DFEDTARL, RRPONTSYAWARD]).
@@ 4.2 The data and the transformations
The remaining parameters $\xi^{k}$ for $k=2,\ldots,K$ are also points on the overnight forward curve, so they tend to move in tandem with the SOFR. For nearer maturities the link tends to be stronger. This suggests modelling the parameters $\xi^{k}$ through the differences
@@
A complete description of the forward curve dynamics requires a model for the lower limit $L$ of the federal funds target range. There are several macroeconomic factors that may affect the FOMC rate decisions, but the most significant ones are the inflation $I$ and the real GDP growth rate $G$; see e.g. [mishkin2007economics]. This is in line also with the Taylor rule [taylor1993discretion], which models the central bank policy rate as a simple parametric function of $G$ and $I$.
We will model the values of $L$, II, and $G$ at FOMC meeting dates. Instead of modelling the values of $L$ directly, we will model the variable $y^{1}:=\ln(L+c^{L})$ as a stationary stochastic process. Here $c^{L}$ is a small positive constant which allows for the possibility of strictly negative target rates in the future. Such scenarios seem plausible given what happened in the Eurozone after the 2008 financial crisis. We choose the value $c^{L}=0.5\%$ which makes $y^{1}$ fairly stationary and symmetrically distributed. We will develop a stationary stochastic model for the three-dimensional process
$y_{t}=[\ln(L_{t}+c^{L}),I_{t},G_{t}].$
@@
Figure 7: Historical values of the macroeconomic risk factors $\ln(L_{t}+0.5\%)$, $I_{t}$, and $G_{t}$ (sources: FRED [DFEDTARL] and Bloomberg Finance L.P.)
@@ 4.3 The macroeconomic model
where $A\in\mathbbm{R}^{3\times 3}$ and $a_{t}\in\mathbbm{R}^{3}$ are given parameters and $\epsilon_{t}$ are iid random vectors in $\mathbbm{R}^{3}$. The above model specification could be generalized in many ways but, as we will see below, the simple specification above already results in a fairly reasonable description of the macroeconomic variables.
@@
We estimate the matrix $A$ using ordinary least squares and set insignificant coefficients to zero. This results in the matrix given in Table 6. The negative diagonal elements indicate a mean-reverting behavior of the components of $y$. The positive signs of $A_{12}$ and $A_{13}$ suggest that the FED tends to raise the target rate $L$ when the inflation $I$ and/or the real GDP growth rate $G$ increase. This is in line with general monetary policy and, in particular, the Taylor rule [taylor1993discretion]. The negative sign of $A_{21}$ reflects the fact that higher interest rates tend to slow down inflation and vice versa. The positive sign of $A_{23}$ means that high real GDP growth tends to increase inflation and vice versa. The negative sign of $A_{32}$ just means that higher inflation tends to slow down real GDP growth.
@@
The above model is close to the macroeconomic model of [ang2003no, Section 4.3], which described the influence of inflation and GDP growth on the interest rate. The model of [ang2003no] did not, however, include the feedback effect from interest rates to the macroeconomy. This effect is an essential feature of the macroeconomy and ce.

### v3

@@ 1 Introduction
Due to the illiquidity and incompleteness, classical arbitrage arguments are not well suited for pricing SOFR derivatives. Indeed, risk-neutral expectations yield linear pricing rules that are poor descriptions of observed market prices or the prices at which an agent might be willing to trade. In the face of incompleteness, the replication arguments behind risk-neutral valuations do not hold. In incomplete and illiquid markets, offered prices are best described by the indifference pricing principle, which is based on approximate hedging arguments capturing nonlinearities and the subjectivity of offered prices. Indifference pricing builds on a subjective (as opposed to risk-neutral) probabilistic description of the underlying rates, prices, and other relevant risk factors such as inflation and GDP; see e.g. [9, 34].
@@
The model presented in this paper captures all the above features. The model includes a description of the development of macroeconomic variables, including the FED rate, on the FOMC meeting dates, which occur roughly on a monthly frequency. SOFR rate and the associated forward curve are then modelled in relation to the FED rate on a daily frequency. This model structure reflects the actual dependence of SOFR on the FED rates, and it captures the jumps observed in SOFR whenever the Fed decides to change the rates. The model is easy to calibrate to historical data, user’s views and observed futures rates. It is fast in simulations, which makes it well-suited for both pricing and risk management applications. We illustrate the simulation model with indifference pricing of SOFR futures, options and swaptions. The pricing exercises here are meant only as an illustration of the convenience of the model in large-scale simulations. The pricing and hedging applications will be developed further in a follow-up paper, that builds on computational optimization of semi-static hedging strategies employing hundreds of exchange-traded SOFR derivatives.
@@
Term structure models under $P$-measure are widely used by public organisations and asset managers. Many of them are based on the Nelson-Siegel and Svensson parameterizations, which describe the yield curve with three or four parameters, respectively; see [29, 40]. For example, [13] develops a stochastic term structure model where the yield curve is statistically connected to macroeconomic variables. An extension with stochastic volatility was proposed in [4] while [43] incorporated corporate bonds. The model presented in this paper allows for more flexible parameterizations, and it incorporates jumps on FOMC meeting dates, which is an essential feature of the overnight lending market.
@@
In derivative pricing, term structure models are often specified so that the corresponding zero-coupon bond prices have arbitrage-free dynamics. A common approach is to specify the term structure dynamics together with a risk-neutral measure (or, more generally, a pricing kernel) under which zero-coupon bond prices of all maturities are martingales. Risk-neutral term structure models for SOFR have been developed in [5, 8, 17, 21, 22, 26, 37, 39] and their references. Statistical analysis in [17] finds that jumps in SOFR are driven by macroeconomic factors, but their stochastic term structure model does not capture the effect. Closer to the present work are the discrete-time models of [1, 24] which describe the joint dynamics of the term structure and relevant macroeconomic variables. In incomplete markets, the martingale measure (or the pricing kernel) is not unique and the chosen one is not necessarily the one that best captures statistical features of the term structure. Like our model, those of [1] and [24] describe the influence of inflation and GDP growth on the central bank policy rate but, contrary to the main role of central banking, the model of [1] does not describe the influence of interest rates on the macroeconomy. Moreover, the models of both [1] and [24] involve latent variables, which makes the calibration of the models complicated. Our model, on the other hand is very easy to calibrate to both historical data, current term structure and user’s views. Neither [1] nor [24] describe the possible jumps on the FOMC meeting dates.
@@ 2.1 SOFR averages and futures
For a three-month SOFR futures contract, the reference period $t_{0},t_{1}]$ starts on the third Wednesday of the third calendar month before the delivery month and ends on the third Wednesday of the delivery month. According to the CME Rulebook [14, Section 46003.A], the payout of a long position is delivered at time $t_{1}$ and given by
@@
where $F(t_{0},t_{1})$ is the futures rate agreed between the counterparties. The exchange-traded three-month futures rates come with bid-ask spreads. We will denote the bid- and ask-rates by $F^{b}(t_{0},t_{1})$ and $F^{a}(t_{0},t_{1})$, respectively. If we want to emphasize that the quotes were observed at time $t$, we will add a subscript and write $F_{t}^{b}(t_{0},t_{1})$ and $F_{t}^{a}(t_{0},t_{1})$.
@@
At CME, SOFR futures are quoted in terms of the “futures price” $f_{t}(t_{0},t_{1})$ which is related to the futures rate above by $F_{t}(t_{0},t_{1})=1-f_{t}(t_{0},t_{1})/100$; see [11].
@@
where $F(t_{0},t_{1})$ is the futures rate agreed between the counterparties.
@@ 3 Parameterization of the forward curve
In practice, the term structure of interest rates is relevant only up to a certain cut-off date $\bar{T}$. For example, if one wishes to estimate the forward curve five years into the future, we would have $\bar{T}=5\times 365$. The number of futures and swap quotes in practice is not enough to uniquely fix the whole forward curve $(F(t))_{t=0}^{\bar{T}}$ through condition (8). Following [20, Section 9.3], we will approximate the overnight forward curve by an affine model of the form
@@
Note that equation (10) is linear in the parameters $\xi=(\xi^{k})_{k=1}^{K}$ so the above least squares problem is quadratic in the parameters $\xi$. Besides futures quotes, one could use quotes on zero-coupon bonds in the calibration. Indeed, zero-coupon bonds imply futures rates $F(t_{0},t_{1})$ through (3) so the corresponding consistency conditions would also be linear. Swap contracts, however, would result in nonlinear consistency conditions for the forward rates so their inclusion in the least squares estimation would result in a non-convex optimisation problem. Such problems are NP-hard, in general; see e.g. [30].
@@
Problem (14) is that of convex quadratic optimization so it can be solved quickly by off-the-shelf optimization solvers. In this paper, we employ the interior point solver of Mosek [2]. The problem was formulated and communicated to Mosek using Python [42] and CVXPY [12]. The solution is obtained in less than 0.4 seconds using a Dell Latitude running Linux with 16 GB of RAM and a 4-core Intel i7 CPU at 3.00 GHz; see Example 4. Similarly, solving 1300 instances of problem (11) takes less than 8 minutes; see Section 4.2.
@@
Such basis functions result in piecewise constant forward curves which are used e.g. by CME [25]. Like CME, we use one-month and three-month futures quotes and SOFR to estimate the forward curve. However, instead of mid-quotes, we use the bid and ask quotes, and solve the least squares problem (14) with tenors corresponding to FOMC meeting dates
@@
$\phi^{k}(t)=\begin{cases}1&t=T_{k},\\ 0&t\notin[T_{k-1},T_{k+1}].\end{cases}$ (15)
This results in continuous piecewise-linear forward curves. Indeed, the corresponding forward curve given by (9), has $F(T_{k})=\xi^{k}$ for all $k$ and the curve interpolates linearly between consecutive tenors. Continuity is often seen as a desirable property in term structure modelling; see e.g. [20, Section 3.3] and the references there. Having an economic interpretation for the parameters $\xi^{k}$, facilitates their statistical modelling; see Section 4.
@@
In particular, the first component of $\xi$ is the one-day maturity zero-rate $F(0)$ which is essentially the current SOFR; see (7). We fit the above parameterization to the observed futures quotes as well as to the value of SOFR observed the morning after. We do this by solving problem (14) with the additional constraint that $\xi^{1}=\ln(1+F(0,1)\delta)/\delta$; see equation (7). This results in the piecewise linear curve in Figure 2. The figure also plots the employed futures bid- and ask-quotes obtained from Bloomberg Finance L.P. Because the SOFR rate is published on the next business day, we shift the horizontal position of the quotes 1 day backwards to match with the futures prices that are published on the trading day.
@@ 4 Statistical modelling of the forward curve
Up to now, we have been studying the term structure at a single point in time. We will next focus on the dynamics of the term structure and denote the forward curve at time $t$ by $F_{t}:=(F_{t}(s))_{s=0}^{T}$. In order to describe the evolution of $F_{t}$, we will, much like in Chapter 9 of [20], first approximate it by the parameterization (9) and then model the parameters as a multivariate stochastic process. More precisely, we will assume that $F_{t}$ has, for every day $t=0,1,2,\ldots$, the parametric representation
$F_{t}(s)=\sum_{k=1}^{K}\xi^{k}_{t}\phi^{k}(s-t)\quad s\geq t,$ (16)
where $\xi^{k}:=(\xi^{k}_{t})_{t=0}^{\infty}$ are real-valued stochastic processes. We will denote by $\xi$ the $\mathbbm{R}^{K}$-valued stochastic process with components $\xi^{k}$. This can be seen as a parametric term structure model in discrete time much as e.g. in [20, Section 9.3]. While term-structure models are usually specified under a fixed risk-neutral measure, we will describe the behavior of the term structure under a “statistical” measure (also known as the $P$-measure” or the “real-world measure”) which will be more suitable for portfolio optimization and indifference pricing of SOFR derivatives. We will follow the general procedure described in [27, Sections 3 and 4] to develop a stochastic model for $\xi$. This allows us to calibrate the model to both historical data and user’s views concerning the future development of the term structure.
@@ 4.1 SOFR and policy rates
SOFR is calculated daily as a volume-weighted median rate of overnight repurchase transactions (repo) of depositary institutions and the FED; see [38]. The FED offers a deposit rate, known as overnight Reverse Repurchase Agreement Facility (ON RRP) since September 2014, and a lending rate, known as Standing Overnight Repurchase Agreement Facility (SRF) since July 2021; see Table 1. Most participants in the repo market have access to these Fed facilities (see [31, 32, 3]) so they are unlikely to lend to a counterparty at a rate lower than ON RRP. Similar to the federal funds target range $[L,H]$, SRF and ON RRP are set during Federal Open Market Committee (FOMC) meetings and remain constant between meetings. Historically, ON RRP was set above $L$, except for a brief period between the end of 2019 and the beginning of 2020; see Figure 3. When enough liquidity exists in the FED reverse repo facility, one would expect SOFR to stay above $L$ as well. The rare cases where SOFR falls below $L$ may be explained by the demand for ON RRP exceeding what the FED is offering, or some participants, which do not have access to FED facilities, accepting deposit rates lower than ON RRP. The converse applies to the federal funds target range upper limit $H$. Table 1 illustrates the situation on 4 January 2024.
@@
Table 1: A glimpse of FED monetary policy measures as of 4 January 2024. SOFR was $5.32\%$ (source: FED [35].)
@@
Figure 3: Daily values of SOFR, ON RRP, and the limits of federal funds target range $[L,H]$ (source: FRED [19, 7, 6, 18]).
@@ 4.2 The data and the transformations
The remaining parameters $\xi^{k}$ for $k=2,\ldots,K$ are also points on the overnight forward curve, so they tend to move in tandem with the SOFR. For nearer maturities the link tends to be stronger. This suggests modelling the parameters $\xi^{k}$, $k=1,\ldots,K$ through the differences
@@
A complete description of the forward curve dynamics requires a model for the lower limit $L$ of the federal funds target range. There are several macroeconomic factors that may affect the FOMC rate decisions, but the most significant ones are the inflation $I$ and the real GDP growth rate $G$; see e.g. [28]. This is in line also with the Taylor rule [41], which models the central bank policy rate as a simple parametric function of $G$ and $I$.
We will model the values of $L$, II, and $G$ at FOMC meeting dates. Instead of modelling the values of $L$ directly, we will model the variable $y^{1}:=\ln(L+c^{L})$ as a stationary stochastic process. Here, $c^{L}$ is a small positive constant which allows for the possibility of strictly negative target rates in the future. Such scenarios seem plausible given what happened in the Eurozone after the 2008 financial crisis. We choose the value $c^{L}=0.5\%$ which makes $y^{1}$ fairly stationary and symmetrically distributed. We will develop a stationary stochastic model for the three-dimensional process
$y_{t}=[\ln(L_{t}+c^{L}),I_{t},G_{t}].$ (19)
@@
Figure 7: Historical values of the macroeconomic risk factors $\ln(L_{t}+0.5\%)$, $I_{t}$, and $G_{t}$ (sources: FRED [6] and Bloomberg Finance L.P.)
@@ 4.3 The macroeconomic model
where $A\in\mathbbm{R}^{3\times 3}$ and $a_{t}\in\mathbbm{R}^{3}$ are given parameters and $\epsilon_{t}$ are iid Gaussian vectors in $\mathbbm{R}^{3}$ with zero mean and covariance matrix $\Sigma$. The above model specification could be generalized in many ways but, as we will see below, the simple specification above already results in a fairly reasonable description of the macroeconomic variables.
@@
We estimate the matrix $A$ using ordinary least squares and set insignificant coefficients to zero. This results in the matrix given in Table 6. The negative diagonal elements indicate a mean-reverting behavior of the components of $y$. The positive signs of $A_{12}$ and $A_{13}$ suggest that the FED tends to raise the target rate $L$ when the inflation $I$ and/or the real GDP growth rate $G$ increase. This is in line with general monetary policy and, in particular, the Taylor rule [41]. The negative sign of $A_{21}$ reflects the fact that higher interest rates tend to slow down inflation and vice versa. The positive sign of $A_{23}$ means that high real GDP growth tends to increase inflation and vice versa. The negative sign of $A_{32}$ just means that higher inflation tends to slow down real GDP growth.
@@
The above model is close to the macroeconomic model of [1, Section 4.3], which described the influence of inflation and GDP growth on the interest rate. The model of [1] did not, however, include the feedback effect from interest rates to the macroeconomy. This effect is an essential feature of the macroeconomy and central banking.